Consider a rock thrown off a bridge of height 79.8
Answer #1

Given data is bridge height h=79.8m,angle\theta=25,initial velocity u=10.5m/s

a)the maximum height h=u2sin2\theta/2g=(10.5)2sin225/2*9.8=1.004m

b)time to reach rock max height t=usin\theta/g=10.5*0.4226/9.8=0.453sec

c)total time to reach land h=-usin\theta+(gt2/2)

79.8=-10.5*sin25*t+(9.8*t2/2)

solve this equation t=4.5sec

d)the horizontal distance cover bythe rock R=u2sin2\theta/g

R=10.52*sin50/9.8=8.6175m

e)final velocity of rock just before reach the ground is v=usin\theta-gt

v=10.5*sin25-9.8*4.5=-39.6626m/s negative direction